OpenStudy (anonymous):

Find the arc length for x(t) = t2 and y(t) = 6t3 − 2 where −1 ≤ t ≤ 1.

OpenStudy (anonymous):

i'm getting the arc length to be zero, but i know that's not right :/

OpenStudy (anonymous):

i'm getting dx/dt to be 2t dy/dt to be 18t^2

OpenStudy (anonymous):

arc length = \[\int\limits_{-1}^{1} \sqrt{4t^2 +324t^4}\]

OpenStudy (anonymous):

= \[\int\limits_{-1}^{1} 2t \sqrt{1+81t^2}\]

OpenStudy (anonymous):

1+81t^2=u

OpenStudy (anonymous):

Du=162t

OpenStudy (anonymous):

\[\frac{ 2(1+81t^2)^{3/2} }{ 243 }\]

OpenStudy (anonymous):

this is what i got when i integrated it without evaluating at -1 and 1

OpenStudy (anonymous):

I hope so

OpenStudy (anonymous):

so is it right?

OpenStudy (anonymous):